A 3.00-kg object has an initial velocity (6i - 2j)m/s. Find the total work done on the object as its velocity changes to (8i + 4j) m/s . (Note: From the definition of the scalar product, v2 = v . v)

Select the correct one:

- Imposible to predict because the displacement its not given

- 72 Joules

-60 Joules

-120 joules
Answer #1

here,

for the initial speed, u = sqrt(6^2 + 2^2) = 6.32 m/s

final speed, v = sqrt(8^2 + 4^2) = 8.94 m/s

for the total work done = change in increase in kinetic energy

the total work done = 0.50 * 3 * (8.94^2 - 6.32^2)

the total work done = 60 J

the total work done is 60 J